# Generalization in Reinforcement Learning for Radio Access Networks

### v1

@@ Abstract
graph neural networks
@@ 3.1 Robust State Reconstruction
RAN functionalities typically operate under partial observability and heterogeneous information aging. Access nodes, for example, can only rely on local or incomplete views of the network state – either from direct sensing or from measurements reported by users or neighboring nodes. Furthermore, such information is inherently affected by uncertainties in its availability when a model needs it. For instance, while a model design may require certain measurements, such data may be temporarily unavailable, outdated, or unsupported depending on the capability of specific network nodes or user devices. Moreover, different observations needed to reconstruct the state of a RAN functionality may be gathered with very diverse timescales. As a result, the state reconstructed at a given time often contains features with different aging rates, leading to incomplete representations of the true state of the underlying system.
@@
This classification allows us to consider different approaches to reconstruct the state of a RAN function. The classic approach is to feed all information, indistinctly, to a deep Q-network (DQN) model. However, to improve generalization across diverse RAN deployments and propagation environments, static and semi-static information can be processed using graph-based methods, such as graph neural networkss Scarselli et al. (2009). Graph structures allow us to embed both radio access nodes and cell attributes, as well as their topological relationships, into rich state representations that can be transferred across different network deployments, improving model robustness to variations in network layout and environmental conditions Suárez-Varela et al. (2023); Shen et al. (2023); Lu et al. (2025). Section 4 further expands on this aspect.
3.2 Training-Environment Diversification
@@ 7.1 Episode design
Figure 4: Markov decision process formulation for downlink link adaptation and HARQ in wireless communication. The figure illustrates how episodes in a MDP are structured to model the interaction between the BS and UE during the downlink transmission and hybrid HARQ protocol. Each episode corresponds to the transmission of a single packet and is composed of a sequence of steps, each defined by a state ($s_{t}$), an action ($a_{t}$), and a reward ($r_{t}$). The process begins with the arrival of a new packet at the BS and proceeds through successive transmissions, receiving NACKs or ACKs from the UE. Colored segments represent repeated transmissions of the same packet due to decoding failures, until successful reception is confirmed via an ACK. This MDP-based abstraction captures the sequential decision-making dynamics inherent in adaptive modulation and retransmission strategies.
@@
As the goal of LA is to adapt the MCS parameters to time- and frequency-varying channel conditions to maintain high link spectral efficiency, we propose a reward design that reflects the spectral efficiency of individual transmissions, as
$r_{n}=SE_{n}\cdot\delta_{n}-\alpha\cdot n\cdot(1-\delta_{n})\qquad n=1,\dots,N,$ (2)
and
$\delta_{n}=\begin{cases}1&\text{if ACK at $n$-th transmission attempt},\\ 0&\text{if NACK at $n$-th transmission attempt},\end{cases}$ (3)
@@
where $n$ denotes the packet transmission counter, the $SE_{n}$ indicates the spectral efficiency of the packet upon transmission attempt $n$, while $\alpha\geq 0$ is a scalar robustness weight used to penalize packet retransmissions.
Optimizing exclusively for spectral efficiency can result in a large number of retransmissions and, thus, lower throughput (which accounts for the time required to deliver a packet). Rule-based LA algorithms like OLLA (cf. Section 6) tackle this by requiring one to adhere to a specific BLER target. Conversely, the reward design in Equation 2 integrates two mechanisms that allow learning policies that discourage exploiting retransmissions to improve spectral efficiency at the expense of throughput. Firstly, upon a successful packet transmission (ACK) at the $n$-th attempt, the value of $SE_{n}$ accounts for the cumulative time-frequency resources used to deliver the data payload with all $n$ transmission attempts. Furthermore, the reward for each failed packet transmission (NACK) is a penalty that increases with each transmission $n$ scaled by $\alpha$.
@@
The reward design Equation 2 is agnostic to the underlying radio transmission technology, scheduling policy, and system bandwidth—a key feature to pursue model generalization over the RAN environment that allow us to deploy a single model in heterogeneous RAN systems with network nodes operating with bandwidth varying from a few to hundreds of MHz. Specifically, using an expression of spectral efficiency renders the reward agnostic the transport block size (TBS) (i.e., information bits) carried by the packet and the number of resource elements $N_{RE}$ allocated by the scheduling policy. Unlike the range of values of TBS and $N_{RE}$, the nominal link spectral efficiency values is agnostic to the system bandwidth and regulated by communication standards such as in the 3GPP NR MCS tables 3rd Generation Partnership Project (3GPP). Furthermore, considering the spectral efficiency per spatial layer further allows us to train a single model that can serve both single-layer and multi-layer MIMO transmissions.
@@ 7.2 Action space design
At its essence, the LA action space consists of a set of MCS index values $\mathcal{A}=\{0,\dots,M-1\}$ reflecting a set of discrete data rates supported by a communication standard. In our evaluations, for instance, we consider the MCS index values defined in the 3GPP 5G NR specifications (e.g., Table 5.1.3.1-1 or Table 5.1.3.1-2 in 3rd Generation Partnership Project (3GPP), supporting modulations orders up to 64QAM and 256QAM, respectively). An action $a_{m}=m$ implicitly provides the modulation order, code rate, and link spectral efficiency to be used for a packet transmission, given a state $s\in\mathcal{S}$. This information, combined with the allocation of time-frequency resource blocks to the UE, enables to determine the TBS used for a new packet transmission according to NR specifications 3rd Generation Partnership Project (3GPP, 3GPP), which can then be used to evaluate the corresponding reward according to Equation 2. Packet re-transmissions, however, reuse the same TBS of the initial packet transmission, while the modulation order and code rate can change.
@@ 7.3 State space design
To enable distributed learning of a single RL policy from the collective experience of all UEs across the network, we model the state of the LA process by combining (a) a UE-centric view of the RAN environment surrounding the UE consisting of semi-static information characterizing network nodes, radio cells and the UE itself (cf. Section 4), and (b) dynamic information characterizing relevant features of the LA process. Together with domain randomization during training (cf. Section 3), this allows us to training models that generalize over the RAN environment.
@@ 7.3.1 Semi-static information
We describe the network deployment surrounding a UE considering its serving cell and subset of most relevant interfering cells, as illustrated in Figure 1, consisting of $K_{I}$ cells whose signal strength measured by the UE exceeds a threshold. As such, the cells comprising the UE’s surrounding may change over time depending on the UE mobility. For each cell, we consider semi-static configuration information, such as the site type (e.g., LTE, NR, etc.), carrier frequency, antenna array configuration (e.g., MIMO or mMIMO), geometry information such as inter-site distance, bandwidth and transmit power configurations, etc.
Furthermore, as the strength and the quality of the signals perceived by a UE can be significantly affected by manufacturing characteristics, such as chipset and antennas hardware and software quality (e.g., algorithms for signal processing and channel state estimation), the MDP state needs to capture information characterizing the UE itself to achieve generalization with respect to the diversity of the UE population. To this end, we consider an indicator representing the UE type and size of the UE antenna array.
In our design, we feed this information either directly to a DQN network or graph-based model (as introduced in Section 4). While a limited set of semi-static parameters suffice to characterize the RAN deployment in our numerical evaluations (cf. Section 8), additional information may be beneficial for practical deployment in live networks.
@@ 7.3.2 Dynamic information
Furthermore, we consider a set of observable information relevant to the LA process of a UE prior to a packet transmission that evolve with fast dynamics and include measurements of the radio link condition, resource availability, and UE state. As LA operates on a sub-millisecond timescale, information aging plays a crucial role to confidently predict the best MCS index to apply for a packet transmission. We therefore apply either a soft- or a hard-aging model to different types of information, as described next.
@@
Packet transmission counter: As LA operates on a per-packet basis, and packet retransmissions may exploit previously received copies of the packet to improve the decoding probability, we include the packet transmission counter $n$ (cf. Equation 2) as part of the state features. The ability to distinguish the first transmission and individual retransmissions helps to learn policies that can optimize LA parameters across multiple transmissions of the same packet.
UE-specific measurements: The 3GPP NR system supports a variety of signals and measurements to observe channel conditions for different purposes. Here, we consider exclusively observations that can be available prior to a packet transmission, such as RSRP measurements and CSI reports. RSRP measurements provide a surrogate to pathloss, which reflects macro channel conditions (e.g., shadowing) and distance between the UE and its serving cell. CSI reports, consisting of a CQI value and a suggested transmission rank, represent the channel dynamics observed by the UE. While the reported CSI typically comply with technical specifications to allow inter-vendor interoperability, differences in UE hardware and proprietary algorithms used to derive CSI reports, can cause UEs of different manufactures to report different CSI observations for the same channel conditions.
Historical CSI: Rather than using a single CSI report, we consider a sequence of $K_{CSI}\geq 1$ historical CQIs and rank values, as well as a weighted average CQI, where the individual CQIs are weighted based on the associated rank.
To account for information aging, we consider a time window $W$ and override CQI and rank values to $-1$ for CSI reports older than $W$. The time window $W$ can be chosen, for instance, based on the channel coherence time, thus allowing to learn that too old information is less relevant to derive LA parameters for the current packet transmission.
@@
UE buffer: The UE buffer status provides insights about data awaiting to be transmitted. Rather than using directly the UE buffer size $B_{t}$ at time $t$, we scale $B_{t}$ by the maximum transport block size $TBS_{\max}$ allowed by the system as
$b_{t}=\min\left\{N_{TTI},\;\frac{B_{t}}{TBS_{\max}}\right\},$ (4)
@@
thereby expressing the buffer size in terms of number of transmissions that would be required to flash the UE buffer, assuming the most favorable channel and resource availability conditions, with $N_{TTI}$ denoting a suitable upper bound. This expression renders information about UE buffer state agnostic to the system bandwidth and supported modulation orders, thus enabling re-usability in heterogeneous RANs.
@@
Expected spectral efficiency (SE): We further consider as part of the LA state an estimate of the link spectral efficiency derived from the most recent historical CSI reports. Specifically, for a given UE $i$, we maintain an average spectral efficiency defined as
@@
where $SE_{k}^{(i)}$ and $rank_{k}^{(i)}$ denotes the link spectral efficiency and the transmission rank, respectively, associated to the $k$-th historical CSI report available for UE $i$. In particular, we model $SE_{k}^{(i)}$ as the nominal spectral efficiency value associated to the $k$-th CQI value reported by the UE as defined, for instance, by the 3GPP 5G NR specifications 3rd Generation Partnership Project (3GPP).
HARQ feedback: Link adaptation algorithms traditionally rely on UE HARQ feedback (i.e., ACK or NACK) to adjust internal link quality estimates for the next transmission. Such information remains relevant also to our design.
We consider a sequence of $K_{H}$ historical HARQ values $\{h_{k}^{(t_{k})}\}_{k=1}^{K_{H}}$, where the superscript $t_{k}$ denotes the time when the $k$-th HARQ value was reported. We model $h_{k}^{(t_{k})}$ as equal 1 if an ACK is reported at time $t_{k}$ and $h_{k}^{(t_{k})}=-1$ for NACK.
However, the relevance of a HARQ value $h_{k}^{(t_{k})}$ to reconstruct the MDP state for a transmission at time $t>t_{k}$ strongly decays with its age $t-t_{k}$. For example, for bursty or ultra-reliable low latency (URLLC) traffic, or in conditions of high cell load, two consecutive transmissions to the same UE may occur with a wide temporal gap (e.g., from tens or hundreds of milliseconds to seconds). As such, a HARQ value $h_{k}^{(t_{k})}$ whose age largely exceeds the channel coherence time may not help reconstruct the current state of the link.
Therefore, rather than using directly the HARQ value $h_{k}^{(t_{k})}$, we define a soft-aging model where $h_{k}^{(t_{k})}$ is linearly scaled based on is age $t-t_{k}$ and a maximum aging $W$ as
$\tilde{h}_{k}^{(t_{k})}=\begin{cases}h_{k}^{(t_{k})}\left(1-\frac{\max\{0,\;t-t_{k}\}}{W}\right)&\text{if}\;t-t_{k}\leq W\\ 0&\text{otherwise,}\end{cases}$ (6)
The soft-aged historical HARQ $\tilde{h}_{k}^{(t_{k})}$ takes values within the interval [1, 1]11[-1,\;1][ - 1 , 1 ], with $\tilde{h}_{k}^{(t_{k})}$ starting from -1 for NACK or 1 for ACK, respectively, and gradually reaching zero as the HARQ age approaches $W$. A value $\tilde{h}_{k}^{(t_{k})}=0$ indicates that the HARQ feedback is too old to provide reliable information for selecting LA parameters for the current transmission.
Historical LA parameters: Link adaptation parameters used in prior transmissions for a UE (e.g., previous MCS index and rank) can provide insights to reconstruct the MDP state for selecting the UE’s LA parameters at time $t$. We therefore consider a sequence of $K_{LA}$ historical MCS index values $\{m_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$ and historical of rank values $\{l_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$. To account for information aging we apply a hard-aging model where historical information older than the maximum age window $W$ is assigned a special value (namely -1) as
$\tilde{m}_{k}^{(t_{k})}=\begin{cases}m_{k}^{(t_{k})}&\text{if}\;t-t_{k}\leq W\\ -1&\text{otherwise.}\end{cases}$ (7)

### v2

@@ Abstract
graph neural network
@@ 3.1 Robust State Reconstruction
RAN functionalities typically operate under partial observability and heterogeneous information aging. Access nodes, for example, can only rely on local or incomplete views of the network state—either from direct sensing or from measurements reported by users or neighboring nodes. Furthermore, such information is inherently affected by uncertainties in its availability when a model needs it. For instance, while a model design may require certain measurements, such data may be temporarily unavailable, outdated, or unsupported depending on the capability of specific network nodes or user devices. Moreover, different observations needed to reconstruct the state of a RAN functionality may be gathered with very diverse timescales. As a result, the state reconstructed at a given time often contains features with different aging rates, leading to incomplete representations of the true state of the underlying system.
@@
This classification allows us to consider different approaches to reconstruct the state of a RAN function. The classic approach is to feed all information, indistinctly, to a deep Q-network (DQN) model. However, to improve generalization across diverse RAN deployments and propagation environments, static and semi-static information can be processed using graph-based methods, such as graph neural networks Scarselli et al. (2009). Graph structures allow us to embed both radio access nodes and cell attributes, as well as their topological relationships, into rich state representations that can be transferred across different network deployments, improving model robustness to variations in network layout and environmental conditions Suárez-Varela et al. (2023); Shen et al. (2023); Lu et al. (2025). Section 4 further expands on this aspect.
3.2 Training Environment Diversification
@@ 7.1 Episode Design
Figure 4: Markov decision process formulation for downlink link adaptation and HARQ in wireless communication. The figure illustrates how episodes in an MDP are structured to model the interaction between the BS and UE during downlink transmission under a hybrid HARQ protocol. Each episode corresponds to the transmission of a single packet and comprises a sequence of steps, each defined by a state ($s_{t}$), an action ($a_{t}$), and a reward ($r_{t}$). The process begins with the arrival of a new packet at the BS (shown in dark blue) and proceeds through repeated transmissions, each followed by either a NACK or an ACK from the UE. Failed transmissions are marked in red, and successful transmissions in green. The corresponding NACK and ACK signals are indicated in light red and light green, respectively. Light blue segments represent retransmissions of the same packet due to decoding failures, repeated until successful reception is confirmed via an ACK.
7.2 Reward Design
@@
Since the goal of LA is to adapt the MCS parameters to channel conditions varying in time and frequency, thereby maintaining high link spectral efficiency, we propose a reward design that directly reflects the spectral efficiency of individual transmissions:
$r_{n}=SE_{n}\cdot\delta_{n}-\alpha\cdot n\cdot(1-\delta_{n}),\quad n\in\{1,\dots,N\},$ (2)
with
$\delta_{n}=\begin{cases}1&\text{if ACK at the $n$-th transmission attempt},\\ 0&\text{if NACK at the $n$-th transmission attempt},\end{cases}$ (3)
@@
where $n$ denotes the packet transmission attempt, $SE_{n}$ indicates the spectral efficiency achieved at attempt $n$, and $\alpha\in\mathbb{R}$ is a scalar robustness weight penalizing retransmissions.
Optimizing exclusively for spectral efficiency may lead to excessive retransmissions, consequently lowering the effective throughput, which considers the total time required to deliver a packet. Rule-based LA algorithms, such as OLLA (see Section 6), address this issue by enforcing adherence to a predefined BLER target. Conversely, the reward formulation in Equation 2 incorporates two complementary mechanisms that encourage policies not to exploit retransmissions excessively, balancing spectral efficiency against throughput. First, when a packet is successfully transmitted (ACK) at the $n$-th attempt, $SE_{n}$ represents the spectral efficiency accounting for the cumulative time-frequency resources utilized across all $n$ transmission attempts. Second, each failed packet transmission (NACK) incurs a penalty proportional to the transmission attempt number $n$, scaled by $\alpha$.
@@
Furthermore, the reward design in Equation 2 is agnostic to the underlying radio transmission technology, scheduling policy, and system bandwidth. This key property facilitates model generalization across heterogeneous RAN environments, enabling the deployment of a single model across nodes operating at bandwidths ranging from a few to hundreds of MHz. Specifically, formulating the reward in terms of spectral efficiency makes it independent of the transport block size (TBS) (i.e., the number of information bits in a packet) and the number of resource elements ($N_{RE}$) allocated by the scheduler. Unlike the wide-ranging values of TBS and $N_{RE}$, spectral efficiency values are standardized and bandwidth-independent, as exemplified by the MCS tables defined in the 3GPP NR specifications 3rd Generation Partnership Project (3GPP). Additionally, measuring spectral efficiency on a per-spatial-layer basis enables the training of a unified model applicable to both single- and multi-layer MIMO transmissions.
@@ 7.3 Action Space Design
The action space of the LA problem consists of a discrete set of MCS indices, represented by $\mathcal{A}=\{0,\dots,M-1\}$, each corresponding to distinct data rates defined by the communication standard. For instance, in our evaluations, we utilize the MCS indices specified by the 3GPP 5G NR specifications (e.g., Tables 5.1.3.1-1 and 5.1.3.1-2 in 3rd Generation Partnership Project (3GPP), which support modulation orders up to 64QAM and 256QAM, respectively).
Given a state $s\in\mathcal{S}$, an action $a_{m}=m$ implicitly determines the modulation order, code rate, and resulting link spectral efficiency for a packet transmission. Together with the allocated time-frequency resource blocks assigned to the UE, this action enables the calculation of the TBS for initial packet transmissions according to NR specifications 3rd Generation Partnership Project (3GPP, 3GPP). The resulting TBS is subsequently used to compute the corresponding reward, as described by Equation 2. For retransmissions, however, the TBS remains unchanged from the initial packet transmission, while the modulation order and code rate may vary.
@@ 7.4 State Space Design
To enable distributed learning of a single RL policy from the collective experience of all UEs across the network, we model the state of the LA process by combining (a) a UE-centric view of the RAN environment surrounding the UE, consisting of semi-static information characterizing network nodes, radio cells, and the UE itself (see Section 4), and (b) dynamic information characterizing relevant features of the LA process. Together with domain randomization during training (see Section 3), this allows us to train models that generalize across the RAN environment.
@@ 7.4.1 Semi-static Information
We describe the network deployment surrounding a UE by considering its serving cell and a subset of the most relevant interfering cells, as illustrated in Figure 1. Specifically, we select $K_{I}$ cells whose signal strength, measured by the UE, exceeds a predefined threshold. Thus, the composition of surrounding cells may change over time depending on UE mobility. For each cell, we include semi-static configuration information such as site type (e.g., LTE or NR), carrier frequency, antenna array configuration (e.g., MIMO or mMIMO), geometric properties (e.g., inter-site distance), bandwidth, and transmit power.
Since the strength and quality of signals perceived by a UE can vary significantly due to manufacturing characteristics—such as chipset, antenna hardware, and algorithms for signal processing or channel state estimation—the MDP state should capture information specific to each UE to generalize effectively across the diversity of the UE population. To this end, we include indicators representing the UE type and antenna array size.
In our design, we provide this information either directly to a DQN or graph-based model, as described in Section 4. While our numerical evaluations (see Section 8) rely on a limited set of semi-static parameters sufficient for characterizing typical RAN deployments, additional parameters may be beneficial for practical deployment in live networks.
@@ 7.4.2 Dynamic Information
We consider dynamic, observable information relevant to the LA process for a UE before packet transmission, encompassing measurements of radio link conditions, resource availability, and the UE’s state. As LA operates on a sub-millisecond timescale, accounting for information aging is crucial to accurately predict the optimal MCS index. We thus apply either a soft or hard aging model to different information types, as detailed below.
@@
Packet Transmission Counter: Since retransmissions can take advantage of previously received copies of the packet to improve the decoding probabilities, we include the packet transmission counter $n$ (see Equation 2) among the state features. Differentiating between initial transmissions and retransmissions allows policies to effectively optimize LA parameters in multiple transmission attempts of the same packet.
UE-Specific Measurements: The 3GPP NR system supports various signals and measurements for channel condition monitoring. We exclusively consider measurements available before packet transmission, including RSRP and CSI reports. RSRP measurements serve as a proxy for pathloss, reflecting macro channel conditions like shadowing and UE-cell distance. CSI reports, comprising a CQI value and recommended transmission rank, capture dynamic channel conditions. While CSI reports adhere to technical specifications facilitating (intervendor) interoperability, hardware and algorithm differences among UEs—made by different manufacturers—may yield variations in reported CSI under identical channel conditions.
Historical CSI: Instead of relying solely on the most recent CSI report, we consider a sequence of $K_{CSI}\geq 1$ historical CQIs and rank values, along with a rank-weighted average CQI. To account for information aging, we define a time window $W$, setting CQI and rank values to $-1$ for reports older than $W$. The window $W$ can correspond to the channel coherence time, allowing models to prioritize recent channel information when determining LA parameters.
@@
UE Buffer Status: The UE’s buffer status indicates the amount of data pending transmission. Instead of using the raw buffer size $B_{t}$ at time $t\in\mathbb{R}$, we normalize $B_{t}$ by the maximum transport block size $TBS_{\max}$ allowed by the system:
$b_{t}=\min\left\{N_{TTI},\;\frac{B_{t}}{TBS_{\max}}\right\}.$ (4)
@@
This expression represents the buffer size as the number of transmissions required to clear the buffer under ideal conditions (i.e., most favorable channel and resource availability), with $N_{TTI}$ serving as an upper bound. This normalization renders the buffer state representation independent of system bandwidth and modulation order, enhancing reusability across heterogeneous RANs.
@@
Expected Spectral Efficiency: We also incorporate into the LA state an estimate of spectral efficiency (SE) derived from recent historical CSI reports. For any given UE $i$, we maintain an average spectral efficiency, defined by
@@
where $SE_{k}^{(i)}$ and $rank_{k}^{(i)}$ denote the spectral efficiency and transmission rank associated with the $k$-th historical CSI report for UE $i$. Specifically, $SE_{k}^{(i)}$ corresponds to the nominal spectral efficiency associated with the $k$-th reported CQI, as defined in the 3GPP 5G NR specifications 3rd Generation Partnership Project (3GPP).
HARQ Feedback: LA algorithms traditionally rely on HARQ feedback (ACK/NACK) from UEs to refine link quality estimates. We consider a sequence of $K_{H}$ historical HARQ values, $\{h_{k}^{(t_{k})}\}_{k=1}^{K_{H}}$, where the superscript $t_{k}$ denotes the time when the $k$-th report was received. We define $h_{k}^{(t_{k})}=1$ for ACK and $h_{k}^{(t_{k})}=-1$ for NACK.
However, the relevance of HARQ feedback diminishes with age—particularly in bursty or ultra-reliable low-latency (URLLC) traffic, or under high cell load, where consecutive transmissions to the same UE may occur with long temporal gaps (e.g., exceeding the channel coherence time). In such cases, outdated HARQ reports may no longer reflect the current link state.
To account for this, we apply a soft-aging model that linearly scales $h_{k}^{(t_{k})}$ based on its age $t-t_{k}$ and a maximum aging window $W$:
$\tilde{h}_{k}^{(t_{k})}=\begin{cases}h_{k}^{(t_{k})}\left(1-\frac{\max\{0,\;t-t_{k}\}}{W}\right)&\text{if }t-t_{k}\leq W\\ \\ 0&\text{otherwise}\end{cases}$ (6)
The resulting value $\tilde{h}_{k}^{(t_{k})}\in[-1,1]$ reflects the aged contribution of each HARQ report, gradually decaying toward zero as its age approaches $W$, thereby reducing its influence on LA decisions when it becomes stale.
Historical LA Parameters: Historical LA parameters (MCS index, rank) from previous transmissions also aid state reconstruction. We include $K_{LA}$ historical MCS indices $\{m_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$ and rank values $\{l_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$. To account for aging, we employ a hard aging model, assigning a value of $-1$ for historical data older than maximum age $W$:
$\tilde{m}_{k}^{(t_{k})}=\begin{cases}m_{k}^{(t_{k})}&\text{if}\;t-t_{k}\leq W\\ -1&\text{otherwise,}\end{cases}$ (7)